OpenStudy (anonymous):

Find the x-intercept and y-intercept of the line 7x-4y=10 is x = -2.5 ? y=1.42 ?

JoelTheBoss (joel_the_boss):

What do you know about this problem?

OpenStudy (anonymous):

you have to substitute y with 0 to find x and x with 0 to find y

JoelTheBoss (joel_the_boss):

Do you know how to do it? Or would you like me to show you?

OpenStudy (godlovesme):

U r correct nice John :)

OpenStudy (godlovesme):

Job***** -_-

OpenStudy (godlovesme):

Wait

OpenStudy (godlovesme):

U switched the answers

JoelTheBoss (joel_the_boss):

To find the x-intercept, substitute in 0 for y and solve for x. \[7x-4(0)=10\]Multiply −4 by 0 to get 0 \[7x+0=10\]Add 7x and 0 to get 7x \[7x=10 \\ x=7/10 \] gtg so hope this helped. Now that you know hat x is you can check for y. :)

OpenStudy (godlovesme):

U mean 10/7 @Joel_the_boss :)

OpenStudy (anonymous):

its 10/7 and -5/2

JoelTheBoss (joel_the_boss):

Yea